2) A monopolist is deciding on the quantity of output to produce in two different countries. Demand for the two countries areNEED HELP PLEASE!
Answer #1

CLASSTIME/Pg. No Date 2 02 = 12-22 24-3p. is man O = 12-P, b) total = 0it02 lunii = 80.333.0 MR-8-0.66667 Mc=4 Profit when MR